OpenStudy (anonymous):

Mr. Klinker is 35 and his daughter is 10. In how many years will Mr. Klinker be twice as old as his daughter?

OpenStudy (lgbasallote):

use this equation \[35 + x = 2(10 + x)\] where x is the number of years later

OpenStudy (anonymous):

\( (35+x)=2(10+x)\) \(35+x=20+2x\) \(15=x\) In 15 years He will be 50 and his daughter will be 25.

OpenStudy (anonymous):

Thank you!